# Instruction Fusion: Advancing Prompt Evolution through Hybridization

### v2

@@ 4 Related Work
Code Large Language Model. The advancement of Large Language Models (LLMs) in open-domain topics has paved the way for extensive research into Code LLMs for code generation tasks. Early studies primarily focused on the pre-training phase of Code LLMs, utilizing models such as Codex Chen et al. (2021b), CodeT5 Wang et al. (2021), StarCoder Li et al. (2023), and Code-Llama. These methods leveraged coding data from open-source platforms like GitHub [^8] https://github.com/ for pre-training. Generating executable and correct codes from these pre-trained LLMs often required intricate prompt engineering. Despite this, such efforts have significantly propelled the progress of code generation tasks and laid a robust foundation for subsequent research in this domain.
Instruction tuning. The reliance on task-specific prompts for extracting information from LLMs, due to its labor-intensive nature and limited generalizability, led to the introduction of instruction tuning Wei et al. (2022). This method enhances the zero-shot capabilities of LLMs in performing tasks via natural language instructions, allowing them to respond to more general human requests. However, relying on human-written instructions or templates limits the quantity, diversity, and creativity of data. In response, Taori et al. (2023) developed Self-Instruct, employing LLMs for both data generation and instruction tuning to create superior synthetic instructions and responses. This technique, utilized by Taori et al. (2023); Roziere et al. (2023), involves collecting high-quality synthetic data from powerful/specialized LLMs for fine-tuning. In the realm of Code LLMs, Code Alpaca Chaudhary (2023) applied Self-Instruct to gather instructions and responses from text-davinci-003999https://platform.openai.com/docs/models/gpt-3-5. Building on synthetic data creation, Xu et al. (2023) proposed Evol-Instruct for evolving instructions to enhance difficulty and diversity. This approach, exemplified in WizardCoder Luo et al. (2023), achieved state-of-the-art performance in Code LLMs but faced evolution process limitations. Our method addresses these issues, advancing code generation performance further. Concurrently, Wei et al. (2023) used LLMs to generate high-quality synthetic data from real-world code snippets, showing promise in code generation. However, as their approach is orthogonal to WizardCoder, it remains similarly orthogonal to ours.
@@ 5 Conclusion
In this paper, we introduce the Instruction Fusion technique, an advancement of Evol-Instruct specifically tailored for code generation tasks. This technique involves the fusion of two evolved instructions into a single, cohesive prompt. It excels in creating instruction sets that are reasonably complex, yet manageable, facilitating a progressive increase in difficulty which is achieved as objectives can be learned separately. The fused instructions also grants a better diversity of the instruction pool. Notably, Large Language Models (LLMs) fine-tuned with this method have demonstrated superior performance across the top five benchmarks, setting new state-of-the-art records.
@@
The primary limitation of the Instruction Fusion method proposed in our study is its cost. We employ GPT-4 Turbo as the Teacher Language Model (LLM), and the fusion prompt incorporates a higher token count compared to the Evol-Instruction. This results in increased expenses for data collection, which cost around 2,200 USD in total. However, it is worth noting compared with manual annotation. Moreover, such costs are rapidly decreasing, thanks to the swift advancements in LLM technology. Another point of concern is the success rate of the fusion process. Fusion of Python-based instructions achieves a pass rate of 93% (where 7% of the fused instructions are deemed unsolvable by GPT-4 Turbo). In contrast, cross-language fusion exhibits a lower pass rate of approximately 63%. This discrepancy arises because many tasks are typically resolved using specific programming languages. Consequently, fusing instructions from different programming languages often leads to impractical outcomes. To address this issue, we propose the development of new prompts that categorize tasks prior to sampling, a solution which could be to explore in future research.

### v4

@@ 4 Related Work
4.1 Code Large Language Model
The advancement of Large Language Models (LLMs) in open-domain topics has paved the way for extensive research into Code LLMs for code generation tasks. Early studies primarily focused on the pre-training phase of Code LLMs, utilizing models such as Codex Chen et al. (2021b), CodeT5 Wang et al. (2021), StarCoder Li et al. (2023), and Code-Llama. These methods leveraged coding data from open-source platforms like GitHub [^8] https://github.com/ for pre-training. Generating executable and correct codes from these pre-trained LLMs often required intricate prompt engineering. Despite this, such efforts have significantly propelled the progress of code generation tasks and laid a robust foundation for subsequent research in this domain.
4.2 Instruction tuning
The reliance on task-specific prompts for extracting information from LLMs, due to its labor-intensive nature and limited generalizability, led to the introduction of instruction tuning Wei et al. (2022). This method enhances the zero-shot capabilities of LLMs in performing tasks via natural language instructions, allowing them to respond to more general human requests. However, relying on human-written instructions or templates limits the quantity, diversity, and creativity of data.
In response, Taori et al. (2023) developed Self-Instruct, employing LLMs for both data generation and instruction tuning to create superior synthetic instructions and responses. This technique, utilized by Taori et al. (2023); Roziere et al. (2023), involves collecting high-quality synthetic data from powerful/specialized LLMs for fine-tuning. In the realm of Code LLMs, Code Alpaca Chaudhary (2023) applied Self-Instruct to gather instructions and responses from text-davinci-003999https://platform.openai.com/docs/models/gpt-3-5. Building on synthetic data creation, Xu et al. (2023) proposed Evol-Instruct for evolving instructions to enhance difficulty and diversity. This approach, exemplified in WizardCoder Luo et al. (2023), achieved state-of-the-art performance in Code LLMs but faced evolution process limitations.
Concurrently, Wei et al. (2023) used LLMs to generate high-quality synthetic data from real-world code snippets, showing promise in code generation. However, as their approach is orthogonal to WizardCoder, it remains similarly orthogonal to ours.
@@ 5 Conclusion
In this paper, we introduce the Instruction Fusion technique, an advancement of Evol-Instruct specifically tailored for code generation tasks. This technique involves the fusion of two evolved instructions into a single, cohesive prompt. It excels in creating instruction sets that are reasonably complex, facilitating a progressive increase in difficulty which is achieved as objectives can be learned separately. The fused instructions also grants a better diversity of the instruction pool. Notably, Large Language Models (LLMs) fine-tuned with IF have demonstrated superior performance across the top five benchmarks, setting new state-of-the-art records.
@@
The primary limitation of the Instruction Fusion method proposed in our study is its cost. We employ GPT-4 Turbo as the Teacher Language Model (LLM), and the fusion prompt incorporates a higher token count compared to the Evol-Instruction. This results in increased expenses for data collection, which cost around 2,200 USD in total for 100k samples (responses included). However, it is worth noting compared with manual annotation. Moreover, such costs are rapidly decreasing, thanks to the swift advancements in LLM technology.
Another point of concern is the success rate of the fusion process. Fusion of Python-based instructions achieves a pass rate of 93% (where 7% of the fused instructions are deemed unsolvable by GPT-4 Turbo). In contrast, cross-language fusion exhibits a lower pass rate of approximately 63%. This discrepancy arises because many tasks are typically resolved using specific programming languages. Consequently, fusing instructions from different programming languages often leads to impractical outcomes. To address this issue, we propose the development of new prompts that categorize tasks prior to sampling, a solution which could be to explore in future research.
Ethics Statement
In our study, we employ GPT-4 for the purpose of amalgamating seed coding instructions, which can not generate sensitive data such as personal information. Our seed instructions are derived from the evol-codealpaca-v1 dataset, a well-acknowledged resource within the open-source community.
Appendix A Instruction Fusion with Magicoder Dataset
In this section, we assess the performance of CodeLlama when employing both our method and the Magicoder method. Specifically, we fine-tune CodeLlama using three datasets: evol-codealpaca-v1, our fused instruction set, and instructions generated in the Magicoder paper.
Table 5: Results of experiments using Magicoder dataset. All models are fine-tuned on CodeLlama. ‘MC’ represents the Magicoder dataset.
[…]
As shown in Table 5, we fine-tune the base models using all three datasets to achieve enhanced performance. Notably, at the time we test the fine-tuned models, our 34B version surpasses all open-source models in the HumanEval and HumanEval+ benchmarks on the EvalPlus leaderboard. However, the lower performance in MBPP and MBPP+ could be attributed to either insufficient prior knowledge, as discussed in Section 3.2, or the use of DeepSeek-Coder, a recently released base model that significantly outperforms CodeLlama, especially on MBPP and MBPP+. For example, the two top-performing models on the leaderboard, which are DeepSeek-Coder-instruct and WizardCoder-V1.1, all utilize DeepSeek-Coder as the base LLM.
[…]